Consider a list of randomly generated 2-letter "words" printed on a paper. The letters cannot be repeated.
(a) At least how many of these "words" should be printed to be sure of having at least 9 identical "words" on the list?
Answer =
(b) At least how many identical "words" are printed if there are 6501 "words" on the list?
Answer =

Answer by sudhanshu_kmr(1152)

maximum no. of different 2 letter words = 26*25 = 650

(a) 650*8 = 5200

5200 + 1 = 5201
to sure 9 identical words 5201 words should be printed.




(b) 6501 = 650*10 + 1

so, there are at least 11 identical words.




If your concept have not been cleared, you are welcome for further help
please feel free to contact.....
sudhanshu.cochin@gmail.com , sudhanshu.cochin@yahoo.com